# Adaptive KDE for Real-Time Thresholding: Prioritized Queues for Financial Crime Investigation

### v1

@@ Abstract
We study the problem of converting a stream of risk scores into one or more review queues under explicit intake constraints[cite: 6]. Instead of top-$K$ or manually tuned cutoffs, we fit an online adaptive kernel density to the score stream, transform the density into a tail-mass curve to meet capacity, and “snap” the resulting cut to a persistent density valley detected across bandwidths[cite: 7]. The procedure is label-free, supports multi-queue routing, and operates in real time with sliding windows or exponential forgetting[cite: 8]. On synthetic, drifting, multimodal streams, the method achieves competitive capacity adherence while reducing threshold jitter[cite: 9]. Updates cost $O(G)$ per event with constant memory per activity.
@@ 1 Introduction
Modern detection systems emit a continuous stream of risk scores. Operations must convert this stream into one or more review queues under explicit intake constraints, while the score distribution itself is non-stationary, multimodal, and often discretized. In practice, this conversion is typically handled by a top-$K$ rule or by hand-tuned cutoffs. Such rules are fragile: small distributional drift can move a fixed cutoff from a sparse region into a dense one, creating abrupt swings in alert volume; adding secondary queues forces ad-hoc adjustments; and repeated manual retuning couples’ analysts’ workload to incidental fluctuations rather than to the underlying structure of the score space.
@@ 3.1 Decision variables and routing
with the understood convention that ”escalation” receives scores in $c_{m},1]$, ”standard” (if present) receives scores $c_{m-1},c_{m})$, and all lower scores hibernate. Within any queue, ordering by score remains monotone (no re-ranking).
@@ 6.2 Two Thresholds (Three Queues)
Figure (5b) depicts two thresholds partition the density into Escalation, Standard, and Hibernation. Boundaries sit in valleys, so small models or traffic wiggles do not flip large numbers of cases across queues. The annotations show expected fractions per queue; within each queue, ordering remains purely by score, enabling staged intake and SLA-aware routing in real time.
@@ 9 Conclusion
We presented a density-anchored, capacity-aware method for prioritized queue construction from streaming risk scores. The approach maintains an online adaptive KDE with boundary correction and uses density valleys as stable, explainable threshold candidates. Capacity is met by snapping cuts to valleys and applying minimal intra-band adjustment.
The policy preserves score monotonicity within queues, supports single or multiple thresholds, and operates in real time without manual tuning. Anchoring cuts at density minima reduces intake sensitivity, yielding smoother queues and portable behavior across cohorts and business activities.
Because the method is nonparametric and score-only, it generalizes across systems and produces an audit trail for threshold placement. Experiments on synthetic streams show reliable capacity matching with substantially lower volatility and cut jitter than quantile baselines.
We do not claim precision gains absent labels. Instead, we provide a transparent, stable, and operationally grounded routing mechanism. Future extensions can add selective label awareness and uncertainty without abandoning the core principle: discover structure first, then route along it.

### v2

@@ Abstract
We study the problem of converting a continuous stream of risk scores into stable decision thresholds under non-stationary score distributions. This problem arises in a wide range of detection systems where scores must be partitioned into prioritized processing regions while preserving semantic consistency over time.
We propose a density-anchored thresholding framework based on online adaptive kernel density estimation. The method summarizes the evolving score population in real time, identifies persistent local minima of the estimated density, and restricts threshold placement to these population-intrinsic separation points. Operational constraints are enforced by selecting thresholds whose induced tail mass satisfies prescribed limits, while anchoring decisions to stable structural features of the score distribution.
The resulting thresholds are label-free, explainable, and robust to distributional drift, discretization effects, and transient fluctuations. The framework naturally extends to multiple thresholds, enabling the construction of ordered processing regions without altering score monotonicity.
Although motivated by financial-crime detection systems, the formulation is generic and applies to any scored stream requiring stable, capacity-constrained routing under evolving score populations.
@@ 1 Introduction
Modern detection systems emit a continuous stream of risk scores. Operations must convert this stream into one or more review queues under explicit intake constraints, while the score distribution itself is non-stationary, multimodal, and often discretized. In practice, this conversion is typically handled by a top-$K$ rule or by hand-tuned cutoffs. Such rules are fragile: small distributional drift can move a fixed cutoff from a sparse region into a dense one, creating abrupt swings in alert volume; adding secondary queues forces ad-hoc adjustments; and repeated manual retuning couples’ analysts’ workload to incidental fluctuations rather than to the underlying structure of the score space. Importantly, the difficulty is not merely operational but structural. The core challenge is to define decision boundaries that remain semantically stable as the score distribution evolves, while respecting externally imposed intake constraints. In this view, capacity does not determine where risk begins or ends; it constrains how much of an already-defined score region can be acted upon at any point in time.
@@ 3.1 Decision variables and routing
with the understood convention that "escalation" receives scores in $c_{m},1]$, "standard" (if present) receives scores $c_{m-1},c_{m})$, and all lower scores hibernate. Within any queue, ordering by score remains monotone (no re-ranking).
@@
Thresholds define fixed score regions whose interpretation is invariant over time. Routing decisions operate only on the volume induced by these regions, not on their semantic meaning. This separation ensures that operational constraints do not implicitly redefine risk levels as workload or traffic fluctuates.
@@ 6.2 Two Thresholds (Three Queues)
Figure (5b) depicts two thresholds partition the density into Escalation, Standard, and Hibernation. Boundaries sit in valleys, so small models or traffic wiggles do not flip large numbers of cases across queues. The annotations show expected fractions per queue; within each queue, ordering remains purely by score, enabling staged intake and SLA-aware routing in real time. The lowest-priority interval produced by this construction naturally serves as a deferred or hibernation region: cases are preserved for future consideration without altering score ordering or boundary semantics. This mechanism applies equally in detection settings where temporary load spikes must be absorbed without distorting score interpretation, including both financial-crime and fraud detection systems.
@@ 9 Conclusion
We presented a density-anchored, capacity-aware method for building prioritized queues from streaming risk scores. The approach treats the score stream as a one-dimensional stochastic object, maintains an online adaptive KDE with boundary correction, and uses the resulting valleys as stable, explainable candidates for thresholds. Capacity is met by snapping cuts to these natural boundaries and, only if needed, making a small intra-band adjustment. The policy preserves monotone ordering by score within each queue, supports single or multiple thresholds for escalation/standard/hibernation, and operates in real time without manual trial-and-error.
Throughout, capacity is treated as a constraint applied to structurally defined score regions, rather than as a driver of threshold placement.
The value is operational and governance centric. Anchoring cuts at density minima reduces sensitivity of intake to small setting changes, yielding predictable volumes, smoother queues, and portable behavior across cohorts and business activities. Because the machinery is nonparametric and score-only, it travels across tenants and systems, producing an audit trail for why thresholds sit where they do. In comparative experiments on synthetic streams, the method matches capacity as reliably as quantile baselines while exhibiting markedly lower volatility and cut jitter, especially under discretization and transient surges.
We do not claim improvements in precision absent outcomes; instead, we provide a transparent mechanism for real-time routing that is robust to incidental fluctuations and easy to justify. Future extensions—band-level yield estimation with uncertainty, density-ratio novelty signals, conditional segmentation with a few operational covariates, and formal queueing-control coupling—can add selective label awareness and broader policy intelligence without abandoning the core principle: discover structure first, then route along it.